# G2L: From Giga-Scale to Cancer-Specific Large-Scale Pathology Foundation Models via Knowledge Distillation

### v1

@@ Introduction
Foundation models (FMs) have been increasingly adopted in computational pathology to extract morphological features from whole-slide images (WSIs). Especially, vision transformers (ViTs) trained with contrastive learning (Chen et al. 2020; He et al. 2020; Caron et al. 2021; Oquab et al. 2023) or masked image modeling (He et al. 2022; Zhou et al. 2022) have achieved remarkable performance across various downstream applications, such as tumor classification (Brancati et al. 2022; Spanhol et al. 2015) and gene mutation detection (Qu et al. 2021).
Recent studies have shown that scaling the amount of training data, diversifying cancer types, and increasing model size consistently improve the performance of FMs (Zimmermann et al. 2024; Filiot et al. 2024). For instance, giga-scale FMs, such as H-optimus-0 (Saillard et al. 2024) and GigaPath (Xu et al. 2024), employ a ViT-Giga (ViT-G) backbone network with 1.9B parameters (Dosovitskiy et al. 2021) and are trained on hundreds of thousands of pathology slides (e.g., 170K slides for GigaPath) that cover a wide range of cancer types (e.g., 28 types for GigaPath). Consequently, the gigascale models outperform their smaller-parameter counterparts, such as ViT-Huge (ViT-H) models with 0.6B and ViT-Large (ViT-L) models with 0.3B parameters, in a variety of downstream tasks.
@@
However, the computational demands of giga-scale FMs, which exceed a billion parameters and require hundreds of thousands of slides for training, present significant challenges for both model development and deployment. In particular, these resource-intensive requirements can limit accessibility in research institutions and clinical settings without a scalable computing infrastructure. In addition, while giga-scale FMs serve as a general feature extractor across various types of cancer, they can possibly dilute critical morphological signals specific to a certain cancer due to their significant heterogeneity (Zhao et al. 2025; Weigelt, Geyer, and Reis-Filho 2010). For example, architectural patterns, nuclear morphology, and stromal interactions unique to breast carcinoma (Weigelt, Geyer, and Reis-Filho 2010) may be underrepresented when training data are dominated by other cancers. As a result, its performance can possibly be degraded for fine-grained discrimination within a specific cancer morphology.
@@
In both breast and prostate cancers, we have evaluated the proposed G2L-distilled model against not only other large-scale FMs but also huge- and giga-scale counterparts across a wide range of downstream tasks, such as gene mutation prediction (Qu et al. 2021), tumor grade classification (Arvaniti et al. 2018), and immune infiltration detection (Saltz et al. 2018; Abousamra et al. 2022). Our results showed that the distilled model outperformed existing large-scale FMs and even surpassed some huge- and giga-scale FMs in a set of benchmarks. These findings demonstrate that the proposed approach is a data- and parameter-efficient method to develop a high-performance, cancer-specific large-scale FM. Fig. 1 compares the G2L-distilled and other FMs based on the TP53 mutation prediction accuracy, model sizes, and inference times, showcasing its superiority over the others.
@@
Fig. 2 illustrates the proposed G2L framework to build a cancer-specific, large-scale FM via knowledge distillation. First, a single cancer type that we want to target for the distillation (i.e., target cancer) is chosen in multi-cancer database (e.g., the cancer genome atlas (TCGA) (Weinstein et al. 2013)). Then, we select 1K pathology slides of the target cancer, where each of these slides is divided into a set of patches in the foreground tissue regions (Lee 2025) (Fig. 2a). Finally, we distilled the knowledge of a giga-scale FM as a teacher to a large-scale FM as a student (Miles and Mikolajczyk 2024) (Fig. 2b).
@@
For giga- and large-scale FMs, H-optimus-0 (Saillard et al. 2024) (ViT-G/14 with 1.9B parameters), which has reported state-of-the-art performance in many downstream tasks in various cancers, was used as a teacher, and Hibou-L (Nechaev, Pchelnikov, and Ivanova 2024) (ViT-L/14 with 0.3B parameters) was utilized as a student. Since these two FMs have different dimensions of output embeddings (1536 for H-optimus-0 and 1024 for Hibou-L), to match them with each other, we simply attached a linear projection layer with batch normalization at the end of the student network, following (Miles and Mikolajczyk 2024). Finally, the loss function (i.e., Log-Sum Loss in Fig. 2) is defined as follows:
@@
We first investigated whether the student model learns a feature space similar to that of the teacher model after applying the G2L framework. To quantify the similarity, we employed centered kernel alignment (CKA) (Kornblith, Shlens, and Hinton 2019), which measures how similarly two models capture spatial information in their feature embeddings by computing the covariance between them and then normalizing the result to produce a similarity score between 0 and 1, where values closer to 1 indicate higher similarity.
@@
Since we aimed to evaluate the performance of FMs as a general feature extractor without fine-tuning, we have adopted the non-training benchmark method as similar to (Cappadona et al. 2024). Concretely, we extracted 50-dimensional image feature vectors using principal component analysis and measured the performance based on the majority vote of the labels (e.g., tumor-infiltrated or not) of the features of k-nearest neighbors (k = 15). For each downstream benchmark, a 10-fold cross-validation was performed, with training and testing data split at an 8:2 ratio. Here, the training data were not used to fine-tune or train the FMs but only to set the labels of neighbors.
For patch-level downstream benchmarks, each patch was fed into the FMs to extract image features, and the performance was measured accordingly. For slide-level and region-of-interest(ROI)-level downstream tasks, non-overlapping patches from each slide or ROI were fed into the FMs to extract a set of features, which were then mean-pooled to generate an aggregated slide-level or ROI-level representation for evaluation. All downstream benchmarks were either binary or multi-label classification tasks, and performance was measured using accuracy.
For comparison, we have benchmarked not only the original teacher (H-optimus-0 (Saillard et al. 2024)) and student (Hibou-L (Nechaev, Pchelnikov, and Ivanova 2024)) models but also other FMs with large- and huge-scales, including UNI (ViT-L with 307M parameters), UNI-v2 (ViT-H with 632M parameters) (Chen et al. 2024), and Virchow (ViT-H with 632M parameters) (Vorontsov et al. 2023).
@@
is a patch-level benchmark dataset extracted from the TCGA cohort (Saltz et al. 2018). It contains 5,245 patches of breast cancer, where each patch is annotated as binary labels (i.e., tumor-infiltrating lymphocytes or not).
@@
is a slide-level benchmark dataset derived from the TCGA breast cancer cohort (Qu et al. 2021). We randomly sampled 511 slides out of a total of 1,133 slides for the benchmark, excluding any slides used for G2L. The binary label for each slide is TP53-mutation or not.
@@
is a patch-level benchmark dataset that comprises 162 breast cancer slides collected at Radboud University Medical Center (Janowczyk and Madabhushi 2016). The total number of patches is 277,524, labeled as IDC-positive or IDC-negative.
@@
is an ROI-level benchmark dataset that contains 4,537 ROIs of 547 pathology slides from 189 breast cancer patients (Brancati et al. 2022). The label for each ROI covers seven subtypes: normal, benign, usual ductal hyperplasia, flat epithelial atypia, atypical ductal hyperplasia, ductal carcinoma in situ, and invasive carcinoma.
@@
is an ROI-level benchmark dataset consisting of 7,909 ROIs from 82 slides (24 benign, 58 malignant) collected at the Universidade Federal do Paraná in Brazil (Spanhol et al. 2015). The labels of benign ROIs include adenosis, fibroadenoma, phyllodes tumor, and tubular adenoma whereas the labels of malignant ROIs were ductal carcinoma, lobular carcinoma, mucinous carcinoma, and papillary carcinoma. The dataset provides each ROI image at four magnifications (40×, 100×, 200×, and 400×), enabling comprehensive evaluation across multiple fields of views.
@@
is a patch-level benchmark dataset that includes 77,364 patches of 331 tissue microarray (TMA) core images collected at the Vancouver Prostate Centre (Arvaniti et al. 2018). For patch extraction, each TMA image was divided into non-overlapping patches with a size of 256 by 256, and each patch was assigned a label according to the majority vote of annotations from six pathologists. The labels for each patch comprise four categories: benign, Gleason grade 3, grade 4, and grade 5.
@@
is a patch-level benchmark dataset that has 156,130 patches from 150 prostatectomy and 53 biopsy slides collected at Singapore’s National University Hospital (NUH) (Huo et al. 2022). As similar as the Gleason benchmark, we assigned a label of each patch based on the majority class of annotations as five categories: benign, stroma, and Gleason grade 3, grade 4, and grade 5.
@@
is a slide-level benchmark dataset that incorporates 190 slides from 95 patients collected from several institutions, including Radboud University Medical Center (Chia et al. 2025). Each slide was labeled as biochemical recurrence positive or negative.
@@
Table 2: Benchmark results of existing foundation models with different scales and proposed G2L-distilled model. The top performance is highlighted in bold, and the second-best is in underline. *: evaluated based on binary labels (benign or malignant) and averaged over multiple magnifications. **: evaluated based on sub-typing labels (e.g., adenosis, fibroadenoma, etc.) and averaged over multiple magnifications.
@@
In addition, we quantified robustness of each model against image variations originating from multiple medical centers using the robustness index, following (de Jong, Marcus, and Teuwen 2025). We randomly sampled an equal number of patches (= 80) in each of five tissue classes (invasive tumor, in-situ tumor, tumor-associated stroma, inflamed stroma, and necrosis) spanning 21 medical centers in the TIGER dataset (Shephard et al. 2022). We first sampled a query patch, selected the top five nearest neighbor patches of the query, and measured the ratio of tissue consistency over center consistency (i.e., number of neighbor patches with the same tissue / number of neighbor patches with the same center). Therefore, larger robustness index (>> 1) represents better discrimination of biologically meaningful features than medical centers. We performed five-fold cross-validation for each FM.
@@ Results
Table 3: Robustness index of foundation models. The best index is highlighted in bold.
Model Robustness Index
H-optimus-0 1.0425 ± 0.01
UNI-v2 1.0756 ± 0.02
Virchow 1.0388 ± 0.02
UNI 1.0338 ± 0.06
Hibou-L 0.9108 ± 0.01
G2L (Ours) 1.0880 ± 0.02
@@
Table 3 presents the robustness index of different FMs. With a score of 1.0880, the G2L-distilled model surpassed both the teacher (1.0425) and all other models. These results support the clinical applicability of the G2L model, indicating its superior ability to prioritize biologically meaningful features over image variations originating from different institutions.

### v2

@@ Introduction
Foundation models (FMs) have been increasingly adopted in computational pathology to extract morphological features from whole-slide images (WSIs). Especially, vision transformers (ViTs) trained with contrastive learning (simclr; moco; dino; dinov2) or masked image modeling (MAE; ibot) have achieved remarkable performance across various downstream applications, such as tumor classification (BRCAS; breakHis) and gene mutation detection (TP53).
Recent studies have shown that scaling the amount of training data, diversifying cancer types, and increasing model size consistently improve the performance of FMs (virchowv2; phikonv2). For instance, giga-scale FMs, such as H-optimus-0 (optimus) and GigaPath (gigapath), employ a ViT-Giga (ViT-G) backbone network with 1.9B parameters (vit) and are trained on hundreds of thousands of pathology slides (e.g., 170K slides for GigaPath) that cover a wide range of cancer types (e.g., 28 types for GigaPath). Consequently, the gigascale models outperform their smaller-parameter counterparts, such as ViT-Huge (ViT-H) models with 0.6B and ViT-Large (ViT-L) models with 0.3B parameters, in a variety of downstream tasks.
@@
However, the computational demands of giga-scale FMs, which exceed a billion parameters and require hundreds of thousands of slides for training, present significant challenges for both model development and deployment. In particular, these resource-intensive requirements can limit accessibility in research institutions and clinical settings without a scalable computing infrastructure. In addition, while giga-scale FMs serve as a general feature extractor across various types of cancer, they can possibly dilute critical morphological signals specific to a certain cancer due to their significant heterogeneity (picture; breat_unique). For example, architectural patterns, nuclear morphology, and stromal interactions unique to breast carcinoma (breat_unique) may be underrepresented when training data are dominated by other cancers. As a result, its performance can possibly be degraded for fine-grained discrimination within a specific cancer morphology.
@@
In both breast and prostate cancers, we have evaluated the proposed G2L-distilled model against not only other large-scale FMs but also huge- and giga-scale counterparts across a wide range of downstream tasks, such as gene mutation prediction (TP53), tumor grade classification (gleason), and immune infiltration detection (tils; tils2). Our results showed that the distilled model outperformed existing large-scale FMs and even surpassed some huge- and giga-scale FMs in a set of benchmarks. These findings demonstrate that the proposed approach is a data- and parameter-efficient method to develop a high-performance, cancer-specific large-scale FM. Fig. 1 compares the G2L-distilled and other FMs based on the TP53 mutation prediction accuracy, model sizes, and inference times, showcasing its superiority over the others.
@@
Fig. 2 illustrates the proposed G2L framework to build a cancer-specific, large-scale FM via knowledge distillation. First, a single cancer type that we want to target for the distillation (i.e., target cancer) is chosen in multi-cancer database (e.g., the cancer genome atlas (TCGA) (TCGA)). Then, we select 1K pathology slides of the target cancer, where each of these slides is divided into a set of patches in the foreground tissue regions (pathology_tile_sampling) (Fig. 2a). Finally, we distilled the knowledge of a giga-scale FM as a teacher to a large-scale FM as a student (KD) (Fig. 2b).
@@
For giga- and large-scale FMs, H-optimus-0 (optimus) (ViT-G/14 with 1.9B parameters), which has reported state-of-the-art performance in many downstream tasks in various cancers, was used as a teacher, and Hibou-L (hibou) (ViT-L/14 with 0.3B parameters) was utilized as a student. Since these two FMs have different dimensions of output embeddings (1536 for H-optimus-0 and 1024 for Hibou-L), to match them with each other, we simply attached a linear projection layer with batch normalization at the end of the student network, following (KD). Finally, the loss function (i.e., Log-Sum Loss in Fig. 2) is defined as follows:
@@
We first investigated whether the student model learns a feature space similar to that of the teacher model after applying the G2L framework. To quantify the similarity, we employed centered kernel alignment (CKA) (CKA), which measures how similarly two models capture spatial information in their feature embeddings by computing the covariance between them and then normalizing the result to produce a similarity score between 0 and 1, where values closer to 1 indicate higher similarity.
@@
To evaluate the performance of FMs as a general feature extractor without fine-tuning, we first adopted the non-training benchmark method similar to (squeezing). Concretely, we extracted 50-dimensional image feature vectors using principal component analysis and measured the performance based on the majority vote of the labels (e.g., tumor-infiltrated or not) of the features of k-nearest neighbors ($k$ = 15). For each downstream benchmark, a 10-fold cross-validation was performed. Here, the training data were not used to fine-tune the FMs but only to assign neighbor labels, thereby enabling a purely non-learnable evaluation of embedding quality.
To complement the non-parametric evaluation, we also performed a linear probing (LP) benchmark (TITAN; SiMLP), assessing the adaptability of the extracted representations of FMs for lightweight supervision. We trained a linear classifier on top of the frozen FM features using a 6:2:2 data split (train:validation:test). While the non-training method evaluates the intrinsic discriminative structure of features, this LP analysis verifies whether representations effectively support downstream tasks with minimal supervision.
For patch-level downstream benchmarks (e.g., TILS dataset below), each patch was fed into the FMs to extract image features, and the performance was measured accordingly. For slide-level and region-of-interest(ROI)-level downstream tasks (e.g., TP53 dataset below), non-overlapping patches from each slide or ROI were fed into the FMs to extract a set of features, which were then mean-pooled to generate an aggregated slide-level or ROI-level representation.
All downstream benchmarks were either binary or multi-label classification tasks. In the non-training evaluation, we reported performance using accuracy. In the linear-probing benchmarks, we measured AUCs for binary tasks and one-vs-rest macro AUCs for multi-label classification.
For comparison, we have benchmarked not only the original teacher (H-optimus-0  (optimus)) and student (Hibou-L  (hibou)) models but also other FMs with large- and huge-scales, including UNI (ViT-L with 307M parameters), UNI-v2 (ViT-H with 632M parameters)  (uni), and Virchow (ViT-H with 632M parameters)  (virchow).
@@
is a patch-level benchmark dataset extracted from the TCGA cohort (tils). It contains 5,245 patches of breast cancer, where each patch is annotated as binary labels (i.e., tumor-infiltrating lymphocytes or not).
@@
is a slide-level benchmark dataset derived from the TCGA breast cancer cohort (TP53). We randomly sampled 511 slides out of a total of 1,133 slides for the benchmark, excluding any slides used for G2L. The binary label for each slide is TP53-mutation or not.
@@
is a patch-level benchmark dataset that comprises 162 breast cancer slides collected at Radboud University Medical Center (IDC). The total number of patches is 277,524, labeled as IDC-positive or IDC-negative.
@@
is an ROI-level benchmark dataset that contains 4,537 ROIs of 547 pathology slides from 189 breast cancer patients (BRCAS). The label for each ROI covers seven subtypes: normal, benign, usual ductal hyperplasia, flat epithelial atypia, atypical ductal hyperplasia, ductal carcinoma in situ, and invasive carcinoma.
@@
is an ROI-level benchmark dataset consisting of 7,909 ROIs from 82 slides (24 benign, 58 malignant) collected at the Universidade Federal do Paraná in Brazil (breakHis). The labels of benign ROIs include adenosis, fibroadenoma, phyllodes tumor, and tubular adenoma whereas the labels of malignant ROIs were ductal carcinoma, lobular carcinoma, mucinous carcinoma, and papillary carcinoma. The dataset provides each ROI image at four magnifications (40×, 100×, 200×, and 400×), enabling comprehensive evaluation across multiple fields of views.
@@
is a patch-level benchmark dataset that includes 77,364 patches of 331 tissue microarray (TMA) core images collected at the Vancouver Prostate Centre (gleason). For patch extraction, each TMA image was divided into non-overlapping patches with a size of 256 by 256, and each patch was assigned a label according to the majority vote of annotations from six pathologists. The labels for each patch comprise four categories: benign, Gleason grade 3, grade 4, and grade 5.
@@
is a patch-level benchmark dataset that has 156,130 patches from 150 prostatectomy and 53 biopsy slides collected at Singapore’s National University Hospital (NUH) (aggc). As similar as the Gleason benchmark, we assigned a label of each patch based on the majority class of annotations as five categories: benign, stroma, and Gleason grade 3, grade 4, and grade 5.
@@
is a slide-level benchmark dataset that incorporates 190 slides from 95 patients collected from several institutions, including Radboud University Medical Center (chimera). Each slide was labeled as biochemical recurrence positive or negative.
@@
Table 2: Benchmark results of existing foundation models of different sizes and the proposed G2L-distilled model. We evaluate accuracy in non-training method (Top) and AUC in linear probing (Bottom). The best performance is highlighted in bold, and the second-best is in underline. *: evaluated based on binary labels (benign or malignant) and averaged over multiple magnifications. **: evaluated based on sub-typing labels (e.g., adenosis, fibroadenoma, etc.) and averaged over multiple magnifications.
@@
Linear Probing (AUC)
TILS Breast 0.9822 ± 0.003 0.9788 ± 0.004 0.9819 ± 0.003 0.9827 ± 0.002 0.9782 ± 0.004 0.9838 ± 0.002
TP53 Breast 0.7603 ± 0.052 0.6795 ± 0.054 0.7138 ± 0.070 0.7085 ± 0.076 0.7317 ± 0.066 0.8046 ± 0.048
IDC Breast 0.9778 ± 0.0008 0.9756 ± 0.001 0.9488 ± 0.002 0.9488 ± 0.002 0.9770 ± 0.0009 0.9796 ± 0.0009
BRCAS Breast 0.8226 ± 0.019 0.8489 ± 0.025 0.7943 ± 0.030 0.7809 ± 0.019 0.8146 ± 0.025 0.8156 ± 0.029
BreakHis (Bin.)* Breast 0.9578 0.9630 0.9471 0.9415 0.9657 0.9626
BreakHis (Sub.)** Breast 0.8436 0.8711 0.7653 0.8033 0.8226 0.8465
Gleason Prostate 0.9846 ± 0.0008 0.9790 ± 0.001 0.9721 ± 0.001 0.9708 ± 0.001 0.9774 ± 0.001 0.9841 ± 0.001
AGGC Prostate 0.9955 ± 0.0001 0.9956 ± 0.0002 0.9939 ± 0.0002 0.9921 ± 0.0002 0.9951 ± 0.0001 0.9958 ± 0.0002
CHIMERA Prostate 0.8451 ± 0.090 0.8653 ± 0.106 0.7980 ± 0.092 0.8173 ± 0.081 0.8572 ± 0.079 0.8520 ± 0.103
@@
In addition, we quantified robustness of each model against image variations originating from multiple medical centers using the robustness index, following (robustness). We randomly sampled an equal number of patches (= 40) in each of five tissue classes (invasive tumor, in-situ tumor, tumor-associated stroma, inflamed stroma, and necrosis) spanning 21 medical centers in the TIGER dataset (TIGER). We first sampled a query patch, selected the top $k$ ($\in$ 3, 5, 10, 20) nearest neighbor patches of the query, and measured the ratio of tissue consistency over center consistency (i.e., number of neighbor patches with the same tissue / number of neighbor patches with the same center). Therefore, larger robustness index (>> 1) represents better discrimination of biologically meaningful features than medical centers. We performed five-fold cross-validation for each FM.
@@ Results
Table 3: Robustness index of foundation models across different $k$ values. The best index is highlighted in bold.
Model $k=3$ $k=5$ $k=10$ $k=20$
H-optimus-0 1.0826 $\pm$ 0.03 1.1890 $\pm$ 0.04 1.3730 $\pm$ 0.07 1.8467 $\pm$ 0.12
UNI-v2 1.0682 $\pm$ 0.03 1.2433 $\pm$ 0.05 1.4113 $\pm$ 0.05 1.8548 $\pm$ 0.10
Virchow 1.0137 $\pm$ 0.04 1.1006 $\pm$ 0.03 1.1351 $\pm$ 0.04 1.2850 $\pm$ 0.07
UNI 1.0701 $\pm$ 0.04 1.1460 $\pm$ 0.02 1.4110 $\pm$ 0.04 1.7811 $\pm$ 0.07
Hibou-L 0.9056 $\pm$ 0.03 0.9905 $\pm$ 0.02 1.0879 $\pm$ 0.01 1.1855 $\pm$ 0.08
G2L (Ours) 1.0891 $\pm$ 0.02 1.3002 $\pm$ 0.10 1.5021 $\pm$ 0.06 2.0316 $\pm$ 0.09
@@
Table 3 presents the robustness index of different FMs across different number of nearest neighbors ($k$ = 3, 5, 10, and 20). The G2L-distilled model consistently surpassed both the teacher and all other models. These results support the clinical applicability of the G2L model, indicating its superior ability to prioritize biologically meaningful features over image variations originating from different institutions.